OpenStudy (anonymous):

Give the value of m and c a). y=2x+3 b).y=4x-2 c).y=2x-5 d). y=1/2x-8 E). 3Y=6X-9 F. 1/2Y=2X+3

OpenStudy (anonymous):

can someone please help me with these problem?

OpenStudy (anonymous):

what is c?

OpenStudy (anonymous):

c is y=2x-5

OpenStudy (anonymous):

whats m?

OpenStudy (anonymous):

Are they saying that m is the slope and c is the constant?

OpenStudy (anonymous):

y = mx + c

OpenStudy (anonymous):

oh sorry i don"t have a figure for m and c. That's how the question reads

OpenStudy (anonymous):

Yup that's what i was thinking surdawi

OpenStudy (anonymous):

y=2x+3 y = mx + c what is m&c in this one

OpenStudy (anonymous):

im used to see it as b lol

OpenStudy (anonymous):

Same but in calculus they say c so i guess its the same thing

OpenStudy (anonymous):

m=slope c=y intercept

OpenStudy (anonymous):

y=2x+3 m=2 c=3